OpenStudy (anonymous):

Two joggers are running around an oval track in opposite directions. One jogger runs around the rack in 56 seconds. They meet every 24 seconds. How many seconds does it take the second jogger to run around the track?

OpenStudy (anonymous):

Ok, so we know they must meet twice for every lap one completes, right? And we know one of the joggers runs a lap in 56 seconds. So that means we can set it up like a mixture problem: If one jogger runs a lap in 56 seconds, and another in x seconds then together they will run it in 2(24) seconds.(twice for per lap). The equation would look like this: 1/56+1/x=1/48 1/x=1/48-1/56 1/x=1/42 1=x/42 42=x So the second jogger must take 42 seconds to run around the track.